OpenStudy (anonymous):

what is the sqrt(7942) i need steps, i divided it by two and got 3971*2

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

I assume you have to evaluate this in surd form? This is a tricky one. Hint: Can you see any properties of 3971? In particular, small(ish?) integers that might be a factor of it?

OpenStudy (anonymous):

wait what? what's surd form?

OpenStudy (anonymous):

In a form with a root, e.g. \[4\sqrt{5}\]

OpenStudy (whpalmer4):

hint: factors of 7942 are 2, 11, and 19

OpenStudy (anonymous):

oh ok, thanks, i can actually 7942 goes down to 722

OpenStudy (anonymous):

oh ok and then divide that by 19 and it goes down to 38

OpenStudy (anonymous):

and then tht goes down to 19 when divided by two

OpenStudy (anonymous):

and then u get 19 sqrt(22)

OpenStudy (whpalmer4):

basically, you just have to try the small prime numbers: 2, 3, 5, 7, 11, 13, 19, 23, 29, etc. and see which ones divide the number without a remainder. yes, \(\sqrt{7942} = 19\sqrt{22}\)

OpenStudy (anonymous):

thanks to all of yall

OpenStudy (anonymous):

One very vaguely useful tidbit to take away from this is 19² = 361 if you didn't know that already :p

OpenStudy (anonymous):

alright, i guess i over looked that one fact